OpenStudy (anonymous):

Which expression is equivalent to the one shown below? (4z)^-3 a. -4z^3 b.-12z^3 c.1/12z^3 d.1/64z^3

OpenStudy (anonymous):

D:\[(4z)^{-3} = 1/(4z)^3 = 1/64z^3\]

OpenStudy (anonymous):

wow that was simple. thank you!